How many Joules of energy are required to change 10 gram of ice at -2 οC to water at 20 οC?
1
Expert's answer
2020-04-30T10:41:41-0400

The total heat required to change 10 gram of ice at "-2^{\circ}\rm C" to water at "20^{\circ}\rm C" is given by


"Q=Q_1+Q_2+Q_3"

"Q=c_{ice}m(0^{\circ}{\rm C}-(-2^{\circ}{\rm C}))+\lambda m+c_{water}m(20^{\circ}{\rm C}-0^{\circ}{\rm C})"

"Q=2100\times 0.01\times 2+334000\times 0.01+4200\times 0.01\times 20"

"=4222\:\rm J."